OpenStudy (anonymous):

What is the slope of a line parallel to the line with equation 5x – 2y = 11?

OpenStudy (jdoe0001):

can you solve that for "y"?

OpenStudy (anonymous):

\[\frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

no i dont know how @jdoe0001

OpenStudy (anonymous):

5x-2y=11 change that to the form y=mx+b (m = slope) first by subtracting 5x from both sides. -2y=-5x+11 now divide each side by -2 y=5/2x+11 because the new line is parallel to this one, it will have the same slope so your answer is 5/2

OpenStudy (jdoe0001):

5x – 2y = 11 < subtract 5x from both sides, then divide both by -2

OpenStudy (anonymous):

k @jdoe001

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

notice homeworkgeek' lines that's how you'd solve for "y"